11. More on the corporate valuation model Aa Aa Acme Corp. is expected to generate a free cash flow (FCF) of $10,175.00 million this year (FCF1$10,175.00 million), and the FCF is expected to grow at a rate of 19.00% over the following two years (FCF2 and FCF). After the third year, however, the FCF is expected to grow at a constant rate of 2.10% per year, which will last forever (FCF4). If Acme Corp.s weighted average cost of capital (WACC) is 6.30%, what is the current total firm value of Acme Corp.? O $388,674.23 million O $382,554.98 million $32,283.31 million O $323,895.19 million Acme Corp.s debt has a market value of $242,921 million, and Acme Corp. has no preferred stock. If Acme Corp. has 375 million shares of common stock outstanding, what is Acme Corp.s estimated intrinsic value per share of common stock? O $237.52 O $647.79 O $215.93 O $214.93Companies that have preferred stock outstanding promise to pay a stated dividend for an infinite period. Preferred stock is treated like a perpetuity if the payments last forever. Preferred stocks are considered to be a hybrid of a stock and a bond. For example, one of the major differences between preferred shares and bonds is that the issuing companies can suspend the payment of their preferred dividends without throwing the company into bankruptcy. However, similar to bonds, preferred stockholders receive a fixed payment-their dividend-before the companys residual earnings are paid out to its common stockholders and, as with common stock, preferred stockholders can benefit from an appreciation in the value of the firms stock securities. Consider the following case of National Petroleum Refiners Corporation (NPR): National Petroleum Refiners Corporation (NPR) pays an annual dividend rate of 8.20% on its preferred stock that currently returns 10.99% and has a par value of $100.00 per share, what is the value of NPRs preferred stock? O $111.92 per share O $100.00 per share $74.61 per share $89.53 per share Suppose that there is high unemployment, which causes interest rates to fall, which in turn pulls the preferred stocks yield to 6.59%. The value of the preferred stock will
Answer #1

Current Value of The the Firm = FCF1/(1+Wacc) + FCF2/(1+WACC)2 + FCF3/(1+WACC)3​​​ + FCF4/((WACC-growth)*(1+WACC)3) = 10175/(1+6.30%) + 10175*(1+19%)/(1+6.30%)2 + 10175*(1+19%)2/(1+6.30%)3 + 10175*(1+19%)2*(1+2.1%)/((1+6.30%)3*(6.30%-2.10%) = 323,895.19

Intrinsic Value of Share = (Value of Firm- Debt)/No of Shares = (323,895.19 - 242,921)/375 = 215.93

Dividend payout = 8.2%*100 = 8.2
Required Rate = Dividend /Preferred Stock = 10.99%
Preferred Stock = 8.2/10.99% = 74.61

If Preferred Stock yield = 6.59%
Preferred Stock = 8.2/6.99% = 117.31

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well